How do you find the vertex of y=x^2+6x+4
Answer by RAY100(1637)

y = x^2 +6x +4
.
if you remember that the x vertex = -b/(2a)
.
x= -(6)/(2 *1) = -3
subst into original eqn to find y
y = (-3)^2 +6(-3) +4
y= +9 -18 +4 = -5
therefore vertex is ( -3,-5)
checking, x=-2, y=4-12+4 = -4
x= -4, y= 16 -24 +4 = -4,,,, ok
.
Alternate solution is to complete the square
y=x^2 +6x +4,,,,remember (6/2)^2 = last term
y-4 = ( x^2 +6x +9 ) -9
y-4+9 = (x+3)^2
(y+5) = (x+3) ^2,,,,,,,,,,form (y-k) = (x-h)^2
.
reading directly from form, (-3,-5)